OpenStudy (anonymous):

Using complete sentences, explain how to graph y = -1/4x + 8 using the slope intercept method? PLEASE HELP ME

ganeshie8 (ganeshie8):

Given line is in slope intercept form y = mx + c slope = m = -1/4 y intercept = c = 8

OpenStudy (anonymous):

Thanks! :)

ganeshie8 (ganeshie8):

we can graph the line in two steps 1) goto y-intercept on y axis (8 units)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):